OpenStudy (anonymous):

When you have (a+b)^2 you can change it into a^2+2ab+b^2 How does it work when I am increasing the exponent? Like (a+b)^3,(a+b)^4..

OpenStudy (akashdeepdeb):

Well, for \((a+b)^n\) expansion particularly, there is a thing called, Pascal's Triangle. There is this thing known as Binomial theorem; we use it for the expansion. eg. \[(a+b)^3 = a^3 + b^3 + 3a^2b + 3ab^2\] \[(a+b)^4 = a^4 + 4a^3b + 6a^2b^2 + 4ab^3 + b^4\]

OpenStudy (akashdeepdeb):

Here is a pattern of their coefficients: 1 1 1 2 1 1 3 3 1 1 4 6 4 1 1 5 10 10 5 1 Can you guess the pattern? ;-)

OpenStudy (anonymous):

Yea, yea, I already heard about that. You know, I am 14 years old and I got really interested into math, I am starting with calculus. Done limits and I was doin functions like f(x)=x^2 so f'(x)=? [ I hope you know what i mean]. And I have seen that pattern like (a+b)^2 = a^2 + 2ab + b^2 so I got bit interested in.. I Will try to get what you said and read more about that. Thank you and have a nice day! :)

geerky42 (geerky42):

You may like to read this: http://www.mathsisfun.com/algebra/binomial-theorem.html

geerky42 (geerky42):

Also, this: http://wdjoyner.com/teach/calc1-sage/html/node106.html Similar pattern.

OpenStudy (anonymous):

I`m trying to understand this, too. So a new operation appears in every bigger power of (a+b). But I can`t find any pattern, how to generate this new opperation...